Determine the equation for the inverse function of y=(x-2)^3+5. Thank you!
Answer by stanbon(75887)

inverse function of y=(x-2)^3+5
Interchange x and y to get:
x=(y-2)^3+5
Solve for y as follows:
(y-2)^3=x-5
Take the cube root to get:
y-2= cube root (x-5)
y=[cube root (x-5)]+2
Cheers,
Stan H,